OpenStudy (anonymous):

Nigel currently has a balance of $668.47 in an account he has held for 15 years. He opened the account with an initial deposit of $497. What is the simple interest rate on the account?

OpenStudy (anonymous):

F=P(1+ni).. F= 668.47, P=497 , n= 15 years.. i=? the equation would be \[\frac{ (\frac{ F }{ P } - 1) }{ n}\] i= 0.023 or 2.3% annually

OpenStudy (anonymous):

forgot to include that equation would be equal to i or the interest rate